OpenStudy (anonymous):

help me please.... really don't know.... find dy/dx at point (1,2) of the implicit function (x^2+y^2)^2=25/4xy^2

OpenStudy (anonymous):

stuck... \[=2(x^2+y^2)(2x+2y.\frac{ dy }{ dx })=\frac{ 25 }{ 4 }.y^2+2y.\frac{ dy }{ dx }.\frac{ 25 }{ 4 }x\]

OpenStudy (anonymous):

what about this ...? but im stucking then.. =.=

OpenStudy (anonymous):

ok ok i'll try now

OpenStudy (anonymous):

You can always just substitute the point and just juggle some numbers to isolate dy/dx

OpenStudy (anonymous):

\[\frac{ (x^2+y^2)^2 }{ xy^2 }=\frac{ 25 }{ 4 }\]

OpenStudy (anonymous):

then expand ?

OpenStudy (anonymous):

and differentiate both side ?

OpenStudy (anonymous):

and then take differentiate both sides

OpenStudy (anonymous):

huh ?? i dont get it . ok this is the original question.. 2)

OpenStudy (anonymous):

really ?? im so sorry

OpenStudy (anonymous):

ok so my abobe steps just now is right ?

OpenStudy (anonymous):

above*

OpenStudy (anonymous):

okkkk wait :)

OpenStudy (anonymous):

What are you doing? That second post looks correct

OpenStudy (anonymous):

But all you have to do at that point is substitute the point (2,1) for x and y then solve for dy/dx

OpenStudy (anonymous):

(1,2) sorry

OpenStudy (anonymous):

What whole thing. You are left with a bunch of numbers and dy/dx standing around. Then you isolate it.

OpenStudy (anonymous):

I mean on the second post where he mentioned he is stuck. All that is needed is to substitute the point (1,2) for x and y and then solve for dy/dx which appears in the equation

OpenStudy (anonymous):

\[2(x^2+y^2)(2x+2y.\frac{ dy }{ dx })-\frac{ 25 }{ 4 }y^2+2y.\frac{ dy }{ dx }.\frac{ 25 }{ 4 }x=0\]

OpenStudy (anonymous):

you mean this ?

OpenStudy (anonymous):

(x^2+y^2)^2=(25/4)xy^2 2(x^2+y^2)(2x+2yy')=(25/4)y^2+(25/4)2xyy' x=1, y=2 2(1^2+2^2)(2*1+2*1+2*2*y')=(25/4)2^2+(25/4)2*1*2*y' Solve for y'

OpenStudy (anonymous):

haha ok ok im wrong. sorry im not good in math -.-

OpenStudy (anonymous):

What is the difference. I can't tell

OpenStudy (anonymous):

What is happening here? I'm confused

OpenStudy (anonymous):

thank you to both of u .. all of these really help me ^^ and btw im bot a boy.

OpenStudy (anonymous):

not*